What is the frequency of a wave with a speed of 28.2 m/s and a wavelength of 12.2 m?
1
Expert's answer
2020-11-30T14:52:38-0500

We can find the frequency of the wave from the wave speed formula:


"v=f\lambda,""f=\dfrac{v}{\lambda}=\dfrac{28.2\ \dfrac{m}{s}}{12.2\ m}=2.3\ Hz."

Answer:

"f=2.3\ Hz."